velocity of a body moving with uniform acceleration of 3m/s2 is changed through 30m/s in certain time. average velocity of body during this time is 30m/s distance covered by it during this time is
1)300m 2)200m 3)400m 4)250m
Ans-(1)300m
 

Grade:11

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Let initial speed be u m/s
let final speed be u+30 m/s
Let time be t

acceleration 3m/(sec*sec)

using formula
V=U + a*T

u+30 = u +3(t)

t=10 sec

now Distance = <speed>*time
= 30*10
= 300 m